OpenStudy (anonymous):

A fan manufacturing company produces 150 fans with a production cost of $18000 and 100 fans with a production cost of $13000. Assume that the relationship between the production cost and number of fans produced is linear. 1 find the equation of the relation ship 2 how much does the production of one fan add to the total production cost? 3 what should be the selling price of the fan in order for the company to earn a profit of $ 19000 on 500 fans 4 find the break event point of the company 5 what should be the fixed cost if company want to earn at least $15300 on 400 fans.?

OpenStudy (aum):

You can think of number of fans as the x-value and the production cost as the y-values. They say the relationship is linear which means it is a straight line. You are given two points on the line: (150, 180000) and (100, 13000). Can you find the slope of the line?

OpenStudy (anonymous):

the slope of the line will be y=500x-57000?

OpenStudy (aum):

Slope \(m = \large \frac{y_2-y_1}{x_2-x_1} = \frac{18000-13000}{150-100} = ?\)

OpenStudy (anonymous):

M= 500

OpenStudy (aum):

How are you getting 500? Can you show the work?

OpenStudy (anonymous):

Sorry, I did wrong in my paper.. is 100.

OpenStudy (aum):

Yes. So y = 100x + b. It passes through the point (100, 13000). Find b.

OpenStudy (anonymous):

y=mx+b 18000=100(150)+b b= 3000

OpenStudy (aum):

Yes. 1) y = 100x + 3000 is the equation.

OpenStudy (aum):

2) how much does the production of one fan add to the total production cost? Slope = rise / run. If the run is 1, then slope = rise. So if the x value, which is the number of fans, increases by 1, the y value, which is the production cost, will increase by slope which is 100. So the production cost will go up by $100 to produce one extra fan.

OpenStudy (aum):

3) 3 what should be the selling price of the fan in order for the company to earn a profit of $ 19000 on 500 fans? First find the production cost for making 500 fans by plugging x = 500 into the equation. Then add $19,000 profit to find the selling cost.

OpenStudy (anonymous):

the selling price will be $34000.

OpenStudy (aum):

I am getting a different number. Show your work.

OpenStudy (anonymous):

y=100(500)+3000 y= 53000 then add 19000 it should be 72000.

OpenStudy (aum):

Yes, $72,000.

OpenStudy (aum):

4) Not sure how to find the break-even point without the revenue function. Is the revenue mentioned anywhere in the problem that you forgot to include?

OpenStudy (anonymous):

the revenue function will 100x r(x)= 100x

OpenStudy (aum):

Break-even point is when the cost function = revenue function Equate the two and solve for x.

OpenStudy (aum):

no. something is not right.

OpenStudy (anonymous):

ya.. because the cost function is 100x+3000

OpenStudy (aum):

where did you get the revenue function from?

OpenStudy (anonymous):

i see my notes.. not from the question.

OpenStudy (aum):

you have to run 4 and 5 by your teacher as some info seems to be missing in the problem you posted.

OpenStudy (anonymous):

ok.. thanks for your help =D

OpenStudy (aum):

you are welcome.